# Lower bounds for Seshadri constants on blow ups of $\mathbb{P}^{2}$

### v1

@@ Abstract.
Let $\pi:X_{r}\rightarrow\mathbb{P}^{2}$ be a blow up of $\mathbb{P}^{2}$ at $r$ distinct points $p_{1},p_{2},\dots,p_{r}$. We study lower bounds for Seshadri constants of ample line bundles on $X_{r}$. First, we consider the case when $p_{1},p_{2},\dots,p_{r}$ are on a curve of degree $d\leq 3$. Then we assume that the points are very general and show that $\varepsilon(X_{r},L,x)\geq 1$ for any ample line bundle $L$ and any $x\in X_{r}$ if the Strong SHGH conjecture is true. We explore the relation between bounded negativity and Seshadri constants and study the Seshadri function on $X_{r}$.
@@ 1. Introduction
If $L$ is submaximal at $x$ then there exists a reduced and irreducible curve $C$ passing through $x$ such that $\varepsilon(X,L,x)=\frac{L\cdot C}{\text{mult}_{x}C}$. Such curves are called Seshadri curves of $L$ at $x$. If $\widetilde{C}$ is the strict transform of a Seshadri curve $C$ on $\widetilde{X}(x)$, then by the Hodge index theorem we have $\widetilde{C}^{2}<0$. A reduced irreducible curve $C$ on $X$ is said to be a $(-m)$ curve if $C^{2}=-m$ and $K_{X}\cdot C=m-2$, where $K_{X}$ is the canonical line bundle of $X$.
@@ 2. Blow ups of points on a curve
Let $p_{1},p_{2},\dots,p_{r}$ be r distinct points lying on a curve $C\subseteq\mathbb{P}^{2}$ of degree $d\leq 3$. Let $X_{r}$ the denote the blow up of $\mathbb{P}^{2}$ at $p_{1},p_{2},\dots,p_{r}$. Then for any ample line bundle $L$ on $X_{r}$ and $x\in X_{r}$, we have
@@ 3. Seshadri constants on Blow ups of $r\leq 8$ distinct points
So $C$ is a singular curve whose strict transform is either a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ or a curve in the linear system $|-K_{\widetilde{X}_{r}(x)}|$. Since $C$ is a singular curve, we must have $d=3$. Hence $\widetilde{C}\in|-K_{\widetilde{X}_{r}(x)}|$. So $p_{1},\dots,p_{r}$ belong to the image of $C$ in $\mathbb{P}^{2}$. As $C$ is irreducible, it is clear that $\text{mult}_{x}C=2$. Since $\varepsilon(X_{8},L,x)<1$, the only possibility is
@@
It is shown in [2, Theorem 1.3] that when $r=6$, a fractional Seshadri constant is achieved. More precisely, $\varepsilon(X_{6},-K_{X_{6}},x)=\frac{3}{2}$ for some $x\in X_{6}$.
@@ 4. General blow ups of $\mathbb{P}^{2}$
    $\widetilde{C}^{2}=-1$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=1$, or
@@
    $\widetilde{C}^{2}=-2$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=0$
@@
$\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}.\widetilde{C}}{2}+1>-2.$
@@
As $\widetilde{C}$ is an effective divisor in a smooth projective surface $\widetilde{X}_{r}(x)$, we have $h^{2}(\widetilde{X}_{r}(x),\widetilde{C})=0$. Since $\widetilde{C}^{2}<0$, $h^{0}(\widetilde{X}_{r}(x),\widetilde{C})=1$. Then
$-h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}}{2}>-2.$
Hence $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})\leq 1.$ We consider the following two cases:
@@
$\varepsilon(X_{r},L,x)\geq 1.$
@@
Let $L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}$ be an ample line bundle on $X_{r}$ and let $\pi_{x}:\widetilde{X}_{r}(x)\to X_{r}$ denote the blow up of $X_{r}$ at $x$. We know that $\widetilde{C}^{2}<0$, where $\widetilde{C}$ the is strict transform of $C$. Let $t=\text{mult}_{x}C$. Since $\varepsilon(X_{r},L,x)<1$ we have $t>1$ and $L\cdot C\leq t-1$. By Lemma 4.1, $20>\widetilde{C}^{2}\geq-$ 2 so $C^{2}=t^{2}-i$, where $i=1$ or $2$. Thus using the Hodge index theorem,
@@
So $(t-1)^{2}\geq t^{2}-i$, which reduces to $t\leq\frac{1+i}{2}$ for $i=1,2$. But this happens only if $t=1$, which is a contradiction. ∎
@@
Clearly Theorem 4.2 implies that if Conjecture 1.3 is true, the above question has a negative answer for $\frac{p}{q}<1$. So it is enough to look at $\frac{p}{q}$ with $p\geq q$. For this, it is enough to look for line bundles with Seshadri constant $\frac{p}{q}$, where $p$ is a prime number.
@@ 5. Bounded Negativity Conjecture and Seshadri Constants
Throughout this paper, we are interested in $\mathbb{P}^{2}$ and blow ups of $\mathbb{P}^{2}$ at finitely many points in $\mathbb{P}^{2}$. It is easy to see that every negative self-intersection curve in a 1-point blow up of $\mathbb{P}^{2}$ is $(-1)$-curve. So Conjecture 5.2 is true for $B(\mathbb{P}^{2})=1$. Let $X_{r}$ be the blow up of $r$ very general points in $\mathbb{P}^{2}$. Then Lemma 4.1 implies that if Conjecture 1.3 is true then Conjecture 5.2 is true for $B(X_{r})=-2$.
@@
This guarantees that for $r\leq 8$, Conjecture 5.2is true with $B(X_{r})=-2$
@@ 6. Submaximality of Line Bundles on General Blow Ups of $\mathbb{P}^{2}$
$V_{x}=\{L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}~{}~{}|~{}~{}\sqrt{L^{2}}<n_{i},e>n_{j}+n_{k}+2n_{l}\text{ for distinct }j,k,l\in\{1,\dots,r\}\}.$

### v2

@@ Abstract.
Let $\pi:X_{r}\rightarrow\mathbb{P}^{2}$ be a blow up of $\mathbb{P}^{2}$ at $r$ distinct points $p_{1},p_{2},\dots,p_{r}$. We study lower bounds for Seshadri constants of ample line bundles on $X_{r}$. First, we consider the case when $p_{1},p_{2},\dots,p_{r}$ are on a curve of degree $d\leq 3$. Then we assume that the points are very general and show that $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$ for any ample line bundle $L$ and any $x\in X_{r}$ if the Strong SHGH conjecture is true. We explore the relation between bounded negativity and Seshadri constants and study the Seshadri function on $X_{r}$.
@@ 1. Introduction
If $L$ is submaximal at $x$ then there exists a reduced and irreducible curve $C$ passing through $x$ such that $\varepsilon(X,L,x)=\frac{L\cdot C}{\text{mult}_{x}C}$. Such curves are called Seshadri curves of $L$ at $x$. If $\widetilde{C}$ is the strict transform of a Seshadri curve $C$ on $\widetilde{X}(x)$, then by the Hodge index theorem we have $\widetilde{C}^{2}<0$. A reduced irreducible curve $C$ on $X$ is said to be a $(-m)$-curve if $C^{2}=-m$ and $K_{X}\cdot C=m-2$, where $K_{X}$ is the canonical line bundle of $X$.
@@ 2. Blow ups of points on a curve
Let $p_{1},p_{2},\dots,p_{r}$ be r distinct points lying on a curve $C\subseteq\mathbb{P}^{2}$ of degree $d\leq 3$. Let $X_{r}$ be the denote the blow up of $\mathbb{P}^{2}$ at $p_{1},p_{2},\dots,p_{r}$. Then for any ample line bundle $L$ on $X_{r}$ and $x\in X_{r}$, we have
@@ 3. Seshadri constants on Blow ups of $r\leq 8$ distinct points
So $C$ is a singular curve whose strict transform is either a fixed component of $|-K_{\widetilde{X}_{r}(x)}|$ or a curve in the linear system $|-K_{\widetilde{X}_{r}(x)}|$. Since $C$ is a singular curve, we must have $d=3$. Hence $\widetilde{C}\in|-K_{\widetilde{X}_{r}(x)}|$. So $p_{1},\dots,p_{8}$ belong to the image of $C$ in $\mathbb{P}^{2}$. As $C$ is irreducible, it is clear that $\text{mult}_{x}C=2$. Since $\varepsilon(X_{8},L,x)<1$, the only possibility is
@@
It is shown in [2, Théorème 1.3] that when $r=6$, a fractional Seshadri constant is achieved. More precisely, $\varepsilon(X_{6},-K_{X_{6}},x)=\frac{3}{2}$ for some $x\in X_{6}$.
@@ 4. General blow ups of $\mathbb{P}^{2}$
    $\widetilde{C}^{2}=-1$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=1$,
@@
    $\widetilde{C}^{2}=-1$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=3$,
  4. (4)
    $\widetilde{C}^{2}=-2$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=0$,
  5. (5)
    $\widetilde{C}^{2}=-2$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=2$,
  6. (6)
    $\widetilde{C}^{2}=-3$ and $K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}=1$.
@@
$\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}.\widetilde{C}}{2}+1>-2.$ (4.1)
@@
As $\widetilde{C}$ is an effective divisor in a smooth projective surface $\widetilde{X}_{r}(x)$, we have $h^{2}(\widetilde{X}_{r}(x),\widetilde{C})=0$. Since $\widetilde{C}^{2}<0$, $h^{0}(\widetilde{X}_{r}(x),\widetilde{C})=1$. Then by (4.1),
$-h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=\frac{\widetilde{C}^{2}-K_{\widetilde{X}_{r}(x)}\cdot\widetilde{C}}{2}>-3.$
Hence $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})\leq 2.$ We consider the following three cases:
@@
  • Case 3:
    $h^{1}(\widetilde{X}_{r}(x),\widetilde{C})=2$.
    Similarly as in above cases, we conclude that types (3) or (5) or (6) can occur
@@
$\varepsilon(X_{r},L,x)\geq\frac{1}{2}.$
Further if $\varepsilon(X_{r},L,X)<1$ then $\varepsilon(X_{r},L,x)=\frac{1}{2}$.
@@
Let $L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}$ be an ample line bundle on $X_{r}$ and let $\pi_{x}:\widetilde{X}_{r}(x)\to X_{r}$ denote the blow up of $X_{r}$ at $x$. We know that $\widetilde{C}^{2}<0$, where $\widetilde{C}$ the is strict transform of $C$. Let $t=\text{mult}_{x}C$. Since $\varepsilon(X_{r},L,x)<1$ we have $t>1$ and $L\cdot C\leq t-1$. By Lemma 4.1, $30>\widetilde{C}^{2}\geq-$ 3 so $C^{2}=t^{2}-i$, where $i=1$ or $2$ or $3$. Thus using the Hodge index theorem,
@@
So $(t-1)^{2}\geq t^{2}-i$, which reduces to $t\leq\frac{1+i}{2}$ for $i=1,2,3$. But this happens only if $t=1$ or $t=2$. As $\varepsilon(X_{r},L,x)<1$ we have $t=2$ and hence $\varepsilon(X_{r},L,x)=\frac{1}{2}$. ∎
@@
Clearly Theorem 4.2 implies that if Conjecture 1.3 is true, the above question has a negative answer for $\frac{p}{q}<\frac{1}{2}$. So it is enough to look at $\frac{p}{q}$ with $2p\geq q$. For this, it is enough to look for line bundles with Seshadri constant $\frac{p}{q}$, where $p$ is a prime number.
@@ 5. Bounded Negativity Conjecture and Seshadri Constants
Throughout this paper, we are interested in $\mathbb{P}^{2}$ and blow ups of $\mathbb{P}^{2}$ at finitely many points in $\mathbb{P}^{2}$. It is easy to see that every negative self-intersection curve in a 1-point blow up of $\mathbb{P}^{2}$ is $(-1)$-curve. So Conjecture 5.2 is true for $B(\mathbb{P}^{2})=1$. Let $X_{r}$ be the blow up of $r$ very general points in $\mathbb{P}^{2}$. Then Lemma 4.1 implies that if Conjecture 1.3 is true then Conjecture 5.2 is true for $B(X_{r})=3$.
@@
This guarantees that for $r\leq 8$, Conjecture 5.2 is true with $B(X_{r})=2$
@@ 6. Submaximality of Line Bundles on General Blow Ups of $\mathbb{P}^{2}$
$V_{x}=\{L=eH-n_{1}E_{1}-\dots-n_{r}E_{r}~{}~{}|~{}~{}\sqrt{L^{2}}<n_{j},e>n_{j}+n_{k}+2n_{l}\text{ for distinct }j,k,l\in\{1,\dots,r\}\}.$